# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

### v1

@@ 7. Workshops with a Proof-of-Concept
While extrinsic modalities have been more extensively studied in training contexts (Spelmezan et al., 2009; Jo et al., 2023b; Huang et al., 2025; Elvitigala et al., 2019), intrinsic feedback remains comparatively underexplored and more difficult to adapt dynamically. We therefore focused our follow-up study on investigating the adaptability of intrinsic feedback in practice, providing finer-grained validation of how design space principles translate into embodied exercise contexts.
Building on the findings of the speed dating study, we conducted a workshop with a proof-of-concept (PoC) to investigate how adaptability in intrinsic feedback could be experienced in practice.
7.1. PoC Design
The PoC system approximated a closed loop of sensing, decision-making, and actuation. Two external cameras captured front and side views of participants’ movements, which were processed on a server for real-time pose estimation and compared against ground-truth postures derived from tutorial videos. The system flagged form deviations and signaled adjustments to strap intensity (Fig. 13).
The digital component of the PoC (Fig. 14) consisted of a user interface and a pose-estimation backend. The interface presented participants with a demonstration of the correct form for each exercise, while the backend was powered by MediaPipe joint tracking for real-time estimation of joint positions and angles. Using both front- and side-view cameras, the system primarily tracked key joints relevant to the exercises, and calculated their current angles relative to the target posture.
@@ 7.1. PoC Design
The PoC suit prototype consisted of a skin-fitting shirt equipped with rigid straps across the chest, shoulders, upper arms, and posterior forearms, complemented by elastic straps on the shoulders. The PoC suit was designed around lateral raise and incline dumbbell fly (Fig. 15), as these exercises together span two key degrees of freedom of the shoulder, one of the most common sites of injury in strength training (Hedrick and Wada, 2008). To account for phase adaptation, each strap was mapped to different phases of movement. Although strap adjustments were performed manually by researchers, this setup simulated how a future motorized system could dynamically modulate intrinsic resistance in real time.
For the incline dumbbell fly, r1 supported the execution phase by guiding the elbow from a folded to a straightened position, and r3 reinforced the completion phase posture by maintaining upright torso alignment. Strap r4 provided feedback on the folding limit of the elbow, preventing unsafe flexion and indicating the appropriate end-point of elbow bend. In parallel, l1 offered progressive resistance throughout the motion, ensuring a smooth and safe trajectory of the arms. In the lateral raise, r2 regulated the arm elevation phase by linking torso and shoulder movement, while r4 again controlled elbow flexion, restricting collapse and guiding the fold to its safe range.
@@ 7.2. Protocol
Survey and Interview: After the PoC trial, participants completed a questionnaire on perceived support, comfort, and appropriateness of feedback. While completing the questionnaire, they were asked to think aloud, and we conducted follow-up questions to further explore their reflections on adaptability, effectiveness, and potential improvements.
@@ 7.3. Results
We analyzed the survey responses, and two researchers reviewed the interview recordings, grouping the findings into themes. Our results further validated the FlexGuard design space, particularly its emphasis on phase-specific movement. Figure 16 summarizes participants’ ratings across survey items (1–7 Likert scale), showing generally positive perceptions of on-body feedback aligned with different phases of movement.
@@ 7.3.1. Phase-Specific Adaptations
7.3.2. Correctness vs. Flexibility
A recurring theme was the balance between enforcing correctness and allowing flexibility. The PoC was rated highly for corrective support (mean = 6.0) but lower for comfort (mean = 4.2). Higher feedback intensity offered stronger prevention against incorrect posture but came at the cost of reduced comfort, as reflected in the lower survey ratings for intensity appropriateness, with some participants noting that it felt too strong. Participants appreciated the rigid feedback that restricted unsafe ROM, but also valued elastic support that permitted controlled deviation. As P5 noted, “If it’s too strict, I feel like I can’t really move.”
However, participants also reported that the feedback from the elastic straps was sometimes too subtle, especially when near the intended plane of motion. This limitation contributed to a lower score on the item about staying on the correct movement path (mean = 4.3), as the elastic straps were not calibrated due to equipment constraints. Nevertheless, participants still felt guided during exercise (mean = 5.3), noting that the elastic straps provided a “kickback” force that pushed their arms back on track when they drifted too high or too low from the recommended plane.
7.3.3. Supporting Injury Prevention
Participants recognized the system’s potential to support injury prevention. Participants rated highly on “The suit helped me avoid unsafe ranges of motion” (mean = 5.7). P2 explained, “This suit stops me before going too far.” Similarly, P4 emphasized the value of strong corrective cues: “The suit stops me right away, which is really great if I’m about to hurt myself.” These results validate the role of rigid straps as hard stops for critical safety.
@@ 8.1. Implications for Sports and Fitness Wearables
By situating injury prevention within a design space framework, FlexGuard advances the conversation on how future sports and fitness wearables can evolve from passive tracking systems to interactive training partners. Our findings illustrate how wearables can move beyond performance monitoring to actively guide movement, prevent injury, and support motor learning. This reframing has broader implications for adjacent domains such as rehabilitation, physical education, and recreational training, where embodied safety and skill acquisition are equally central. In this way, FlexGuard demonstrates how design spaces can serve as actionable frameworks for envisioning the next generation of intelligent, interactive wearables.
@@ 8.2. Intelligence in Materials
Our findings show that there is a need for greater intelligence in materials themselves. Many participant needs pointed toward adaptive intrinsic feedback, where feedback is conveyed through material properties such as rigidity or elasticity. Unlike extrinsic actuation, which can be modulated digitally or mechanically, intrinsic feedback relies on the physical response of the material. To make such feedback adaptive, materials themselves must become “smart” and dynamically tunable. This suggests a path forward through mechano- or physical intelligence (Wang and Li, 2023), where material properties are engineered to support highly adaptive, wearable use cases.
@@ 9. Limitations and Future Work
Our work carries several limitations that inform directions for future research. First, we investigated only upper-body movements. While these capture many common injury-prone patterns, they represent only a subset of the body. Some insights may generalize to other regions with similar roles, such as joints that govern ROM, supporting muscle groups, and related structures, but future work should extend our approach to lower-body and full-body exercises, where different biomechanical demands may surface new requirements for sensing and feedback design.
Second, in our PoC validation we relied on manual configuration of straps to set feedback intensity levels. While sufficient for this study, this approach introduced setup delays and required facilitator involvement. Although strength training movements are relatively straightforward, minimizing delays will be especially important when studying on-body feedback in complex, time-sensitive movements. It also remains underexplored how multiple feedback cues delivered in a short time window can be coordinated to minimize cognitive load.
Third, our strap tension measurements (7–15 N across different locations) revealed an important limitation for wearable systems relying on passive, strap-based feedback. Stronger corrective cues require higher strap tensions, yet higher tensions make it more difficult to maintain comfort and stability without bulky motors or anchoring structures. Designing garments that can distribute forces evenly across the body while remaining flexible and wearable in dynamic exercise contexts remains an open challenge.
@@ 10. Conclusion
This paper presented FlexGuard, a design space for on-body wearable feedback for injury prevention in strength training derived from nine co-design workshops with trainer–trainee pairs using a technology probe that explored both extrinsic and intrinsic modalities. FlexGuard is articulated through dimensions of Sensing & Triggering and Feedback Delivery, providing designers with a framework to associate injury risks in specific muscles with appropriate forms of feedback to prevent injuries during exercise. We validated this design space through a speed dating study with seven beginners, which highlighted the importance of designing semantics when using extrinsic modalities, the timing of feedback delivery, and adaptation across fast-changing training contexts. We further validated FlexGuard at the movement level through a PoC prototype evaluated with six beginners, where results revealed tensions between effectiveness and comfort, variability in perception across users, and the challenges of achieving dynamically adaptive feedback. Overall, our findings show that FlexGuard is a valid and useful framework for designing on-body feedback systems aimed at injury prevention in strength training.

### v2

@@ 7. WORKSHOPS WITH A PROOF-OF-CONCEPT
We sought to translate FlexGuard into a proof-of-concept (PoC) prototype and utilize it during training to further validate how the design space operates in practice. In addition, we also aimed to gather additional evidence on users’ underlying needs, several of which were surfaced during the speed dating study. In this section, we outline the scope of this validation, the rationale behind the PoC design, and the study protocol and results.
7.1. Motivation for PoC
Given the breadth of FlexGuard, we focused our scope on Feedback Delivery, as it is the aspect that most requires direct user experience to evaluate in context. While prior work in physical training and related domains has examined portions of this design space, primarily through extrinsic feedback such as vibrotactile systems for snowboarding (Spelmezan et al., 2009) and rock climbing (Kosmalla et al., 2016) (e.g., extrinsic (discrete), instant or delayed alerts), high-fidelity full-body suits like the Teslasuit (Caserman et al., 2021) (e.g., continuous, instant cues), and pneumatic garments such as VabricBeads (Pardomuan et al., 2024) (e.g., extrinsic (continuous), compressive actuation), these systems have largely focused on delivering feedback rather than examining how feedback scaffolds trainees’ experience of safety during movement. While such systems deliver useful support, intrinsic feedback remains comparatively underexplored and yet offers distinct benefits for personalized, intelligent wearables that scaffold safety: it provides continuous, instant cues generated directly through the material properties of the wearable itself, allowing users to remain aware of safety-relevant movement while experiencing embodied support as part of their motion rather than as an external interruption.
Guided by these observations and insights from our previous study, we scoped the PoC to both validate design dimensions through an embodied instantiation of intrinsic feedback and clarify user-needs assumptions surfaced during speed dating. Because our focus is on Feedback Delivery, the PoC operationalizes a set of dimensions that extends beyond what prior work has addressed, including Feedback Type (intrinsic via both rigid and elastic mechanisms), Feedback Intensity (low, high, progressive), Feedback Latency (instant), and Feedback Direction (opposing, guiding), with Feedback Location at leading movement surfaces across the upper body.
7.2. PoC Suit Design
@@
The PoC suit prototype consisted of a skin-fitting shirt equipped with rigid straps across the chest, shoulders, upper arms, and posterior forearms, complemented by elastic straps on the shoulders. The PoC suit was designed around lateral raise and incline dumbbell fly (Figure 13), as these exercises together span two key degrees of freedom of the shoulder, a region associated with common safety concerns and strain in strength training (Hedrick and Wada, 2008). To account for phase adaptation, each strap was mapped to different phases of movement. Although strap adjustments were performed manually by researchers, this setup simulated how a future motorized system could dynamically modulate intrinsic resistance in real time.
Rigid straps (r1–r4) functioned as opposing feedback, constraining movement in undesired directions and offering adjustable low and high intensity levels through strap tightening. In contrast, the elastic strap (l1) acted as guiding feedback, nudging the arm toward the intended posture with progressive intensity, where larger deviations resulted in stronger resistance.
For the incline dumbbell fly, r1 constrained elbow extension as the arms transitioned from folded to straightened positions, while r3 supported upright torso alignment during the hold phase. Strap r4 applied opposing feedback near the folding limit of the elbows, making excessive flexion difficult and indicating a boundary for elbow movement. In parallel, l1 offered guiding progressive resistance throughout the movement phases, helping participants remain aware of arm trajectory within the intended movement plane relative to the body.
In the lateral raise, r2 regulated arm elevation by linking the torso and shoulder, while r4 again constrained elbow flexion, limiting collapse and signaling acceptable movement ranges.
7.3. System Operation
The PoC system, as shown in Figure 14, operated through three phases: sensing, decision-making, and actuation. For sensing, two external cameras captured Front and Side views of participants’ movements, which were streamed to the PoC’s digital interface displayed on a monitor visible to the researchers. These video feeds were processed with MediaPipe (Dill et al., 2023) for real-time pose estimation. Using key postures for each exercise (see Supplementary Material), joint-angle comparisons were computed against ground-truth angles derived from tutorial videos, using the joint landmarks obtained from each camera view. When a real-time joint angle exceeded the allowable threshold in either view, the interface displayed the corresponding angle label in red; when the angle remained within bounds, the label appeared in green (Figure 14’s before-after).
Decision-making was driven by a system notification, illustrated in the notification panel in figure 14, computed from the magnitude and consistency of deviations from the ground-truth posture. For each set of five repetitions, the system calculated the deviation angle for every rep and considered a deviation meaningful only if it exceeded the threshold in a majority of repetitions (i.e., at least 3 out of 5). When this occurred, the system averaged the deviation angles across the above-threshold repetitions to determine the recommended feedback level for that joint. Small averaged deviations (e.g., –) triggered a recommendation for low-intensity feedback, whereas larger averaged deviations (e.g., or more) triggered high-intensity feedback. For motion-path deviations, the system applied an additional buffer to account for reduced MediaPipe performance in the Side view. As before, if more than half of the repetitions exceeded this adjusted threshold, the system recommended progressive feedback via the elastic strap. All of the information will be summarized in a notification after each set of exercise. When deviations did not meet the criteria, the system recommended no feedback for that joint.
We approximated the actuation phase by having researchers manually tighten the straps to simulate adaptive intrinsic feedback. Two researchers continuously monitored the interface, and after each set, the monitor displayed a notification indicating where strap adjustments were needed. Researchers then manually configured the feedback level on the relevant joint while the participant was idle in their rest posture between sets, as illustrated in the before and after comparison panel in Figure 14. We adopted this approach because fully automating intrinsic feedback actuation in real time, while keeping the system wearable and non-intrusive for strength training, is non-trivial and beyond the scope of this paper’s contribution.
@@ 7.4. Protocol
Survey and Interview: After the PoC trial, participants completed a questionnaire on perceived support, comfort, and appropriateness of feedback. While completing the questionnaire, they were asked to think aloud, and we conducted follow-up questions to further explore their reflections on adaptability, effectiveness, and potential improvements. During this phase, participants were also free to retry exercises or explore feedback configurations on the suit that they had not experienced earlier in the session.
@@ 7.5. Results
We analyzed the survey responses, and two researchers reviewed the interview recordings. This analysis focused on evaluating how the PoC validated the proposed design space dimensions outlined in the study scope. We also report recurring themes derived from the study data, some of which were previously identified in the speed dating study. More details about the PoC’s corrective effectiveness can be found from Supplementary Material.
7.5.1. Overview of Evaluated Dimensions
Each of the dimensions was validated through subjective data or objective data from the pose tracking system.
Feedback Type. The PoC evaluated two intrinsic mechanisms, rigid and elastic straps, to examine how participants perceived differences in constraint and flexibility during movement. Several participants described rigid feedback as a strong constraint that helped them avoid unsafe positions: one noted that the straps “helped restrict [me] to overdoing some position because the tension is kind of really tight and strong,” (P2), while another explained that they are “always afraid of overextending” and that the suit “definitely helped restrict me from doing that” (P3). At the same time, others criticized rigid straps as “too restrictive” or uncomfortable when bending the arm. Overall, 4 out of 6 participants (P1-P3, P5) described the rigid straps to limit unsafe motion. In contrast, participants characterized elastic feedback as more supportive than restrictive. One participant stated, “elastic is better because you can still get the feedback but it’s less… I don’t think I need to be restricted a ton, I just need a little bit of feedback and then I can adjust the movement myself,” and another preferred elastic straps because they “give you a pathway back to ground… if you pull, you feel that pull… from both angles” (P4). Overall, four participants expressed a preference for elastic configurations for at least some phases of movement.
Participants generally described rigid feedback as providing a “firm stop” that prevented unsafe motion (P2, P4), while elastic feedback offered “helpful guidance” that supported a natural trajectory without feeling restrictive (P3, P6).
Feedback Intensity. The PoC implemented three levels of Feedback Intensity, i.e., low, high, and progressive, to reflect varying degrees of corrective resistance. Participants described low intensity as a subtle prompt that let them adjust their own form, noting that they “just need a little bit of feedback” (P2) rather than being strongly restricted. high intensity was associated with strong stopping forces that “helped restrict [me] from overdoing some positions,” but several people also found the highest level “too restrictive” (P3) or “more uncomfortable” (P5) and asked either for more resistance in some cases or less in others. progressive intensity was experienced as feedback that increases gradually where one participant explained that with the bending strap “the feedback is from weak to strong,” and others said that they could feel the tension increase when they moved further out of range, which aligns with the survey data (Figure 15) that prefers adjustable or progressive resistance over a single fixed intensity level.
Feedback Latency. The system simulated instant feedback by manually tightening or relaxing straps immediately upon detecting form deviations. Participants highlighted the importance of timing in feedback delivery, describing immediate correction as essential to maintaining safe technique, P1 noted that the straps gave “strong feedback” that prevented them from “going the other way around” or “moving up over the final position,” and that they could “feel feedback appropriately aligned with the different phases of movement.” At the same time, several participants implicitly expressed preferences with delayed feedback, explaining that the suit often corrected posture at the boundaries of the motion rather than during the movement itself. For example, one participant explained, “it helped me fix my posture at the end,” (P2) and another remarked that they “didn’t feel that much feedback during my movements” (P6) and only noticed it “when I’m out of range.” These reflections indicate openness to delayed or boundary-based cues, especially when continuous correction felt unnecessary or overly restrictive.
Feedback Direction. Two directional strategies were tested through the PoC straps: opposing and guiding feedback. Five participants interpreted directional cues as meaningful signals to adjust their movement. opposing feedback provided resistance against incorrect motion paths, while guiding feedback assisted users back toward the correct trajectory. opposing feedback was noted by three participants as resistance that helped push them back if they drift off-path. P5 noted ”I can feel the tension if I’m in the wrong position” and P4 described it as ”You can feel the pull from both angles.” In contrast, guiding feedback was mentioned in four sessions, helping participants understand where their arms should remain without fully restrict the motion. ”It helped [me] restrict from overextending.” (P3), and P6 reported that the elastic strap ”the stretch helped me remain in the correct position instead of like just lying down.”
Together, these findings demonstrate that the PoC successfully embodied the adaptive principles of the FlexGuard design space, validating user needs for responsive, phase-aware, and semantically interpretable on-body feedback mechanisms in safety-oriented training.
@@ 7.5.2. Phase-Specific Adaptations
7.5.3. Movement Continuity in Embodied Use
Consistent with the movement continuity considerations surfaced in the speed dating study, the PoC revealed a clear embodied tension between maintaining movement continuity and applying corrective constraint. The PoC was rated highly for corrective support (mean = 6.0) but lower for comfort (mean = 4.2). Higher levels of constraint (rigid feedback) were experienced as effective for restricting unsafe ranges of motion, but often at the cost of disrupting movement continuity. As P5 noted, “If it’s too strict, I feel like I can’t really move.”
In contrast, participants reported that elastic straps allowed them to continue moving while remaining aware of deviations, enabling adjustment without full interruption. However, this flexibility also introduced limitations. Several participants noted that elastic feedback could be too subtle near the intended plane of motion, contributing to a lower score on staying on the correct movement path (mean = 4.3). This effect was partly due to calibration constraints in the prototype. Despite this, participants still reported feeling guided during exercise (mean = 5.3), describing elastic feedback as providing a “kickback” force that nudged movement back on track rather than stopping it outright. The finding demonstrates how different feedback strategies shift the balance between allowing movement to continue and interrupting it when safety risks increase.
7.5.4. Scaffolding Safety-Relevant Movement
Participants recognized the system’s potential to scaffold safety by supporting their awareness of physical limits and providing embodied constraint at safety-critical moments. Participants rated highly on “The suit helped me avoid unsafe ranges of motion” (mean = 5.7). P2 explained, “This suit stops me before going too far.” Similarly, P4 emphasized the value of strong corrective cues: “The suit stops me right away, which is really great if I’m about to hurt myself.” Together, these results validate the role of rigid straps as movement constraints that scaffold trainees’ awareness of and response to safety-relevant movement at critical moments.
@@ 8.1. Implications for Sports and Fitness Wearables
Across our series of studies, we synthesized a set of design guidelines for future on-body feedback wearables that scaffold safety during strength training. These guidelines highlight how different forms of feedback function in practice and how they can be strategically combined to help users recognize safety-relevant movement while also providing embodied support during training.
8.1.1. Complementary Roles of Intrinsic and Extrinsic Feedback
Our findings show that intrinsic and extrinsic feedback modalities offer distinct strengths, challenges, and experiential qualities in how they scaffold safety during strength training. Table 3 summarizes these differences. These complementary qualities point toward hybrid systems that combine the immediacy and physical grounding of intrinsic feedback with the flexibility and interpretability of extrinsic cues to support trainees’ awareness of and response to safety-relevant movement.
| Dimension | Intrinsic Feedback | Extrinsic Feedback |
| User Experience | • Embodied guidance that stabilizes posture • Clear physical boundaries and ROM cues • May feel restrictive or uncomfortable at higher tensions | • Lightweight cues that do not restrict movement • Effective for timing, attention, and symbolic signaling • Risk of ambiguity if cue semantics are inconsistent |
| Technical Challenges | • Limited adjustability once configured due to reliance on material properties • Physical bulk or tension may impact comfort or mobility • More difficult to encode nuanced or varied feedback meanings | • Requires power, electronics, and careful placement • Harder to convey physical boundaries without relying on user interpretation • Timing and coordination must be carefully calibrated to avoid overload |
| Best-Suited Contexts | • Constraining ROM boundaries • Guiding slow or controlled movement • Providing stability during load-bearing tasks | • Signaling imbalance, tempo deviations, or fatigue • Supporting early-stage learning or technique awareness • Event-driven alerts and motivational or attentional cues |
8.1.2. User Needs for FlexGuard’s Application
While FlexGuard provides a structured way to reason about sensing and on-body feedback, applying the design space in practice requires attention to several user-centered needs. Below, we summarize the key considerations that emerged across our studies.
Adaptability to context. Wearable feedback should flexibly adjust to both rapid changes, such as transitions across movement phases, and slower-changing factors such as training experience, fatigue, or individual body mechanics. This adaptability may involve modulating feedback modality and intensity based on the moment-to-moment demands of the exercise. Subtle cues tended to be helpful during continuous motion, whereas stronger, more restrictive cues were more appropriate near ROM limits or under higher loads. These insights point toward context-aware systems that can modulate feedback in real time while still accommodating longer-term user characteristics.
Interpretable feedback semantics. Feedback should convey clear and consistent meaning to remain actionable. Some signals, such as rigid stops at ROM boundaries, naturally communicate their intent. However, physical constraints alone may not always indicate how to correct posture, guiding users closer to (but not precisely into) the intended position. Extrinsic modalities such as vibrotactile cues can express more nuanced information, though they also rely on how users interpret the signal. When users can understand what a cue means and how to adjust accordingly, feedback becomes a stronger support for learning safe movement patterns. Ensuring that feedback remains understandable is therefore an important design consideration, especially as guidance becomes more complex.
Movement autonomy. Effective systems should support freedom of movement when it is safe while offering firmer constraints when necessary. Excessive restriction during dynamic phases can feel uncomfortable or disrupt natural exercise rhythm, whereas carefully timed constraints at higher-risk points can help prevent unsafe ROM or posture breakdown. These observations suggest the value of systems that intervene selectively, preserving autonomy during safe motion while offering support when risk increases.
8.1.3. Toward Interactive Training Partner
By situating safety scaffolding within a design space framework, FlexGuard advances the conversation on how future sports and fitness wearables can evolve from passive tracking systems to interactive training partners. Our findings illustrate how wearables can move beyond performance monitoring to actively guide movement and support users in recognizing safety-relevant deviations during exercise, while still supporting motor learning. The idea has broader implications for adjacent domains such as rehabilitation, physical education, and recreational training, where embodied safety and skill acquisition are equally central. In this way, FlexGuard demonstrates how design spaces can serve as actionable frameworks for envisioning the next generation of intelligent, interactive wearables.
@@ 8.2. Intelligence in Materials
Our findings show that there is a need for greater intelligence in materials themselves. Many participant needs pointed toward adaptive intrinsic feedback, where feedback is conveyed through material properties such as rigidity or elasticity. Unlike extrinsic actuation, which can be modulated digitally or mechanically, intrinsic feedback relies on the physical response of the material. To make such feedback adaptive, materials themselves must become “smart” and dynamically tunable. This suggests a path forward through mechano- or physical intelligence (Alu et al., 2025), where material properties are engineered to support highly adaptive, wearable use cases.
@@ 9. LIMITATIONS AND FUTURE WORK
Our work carries several limitations that inform directions for future research. First, we investigated only upper-body movements. While these capture many common safety-sensitive patterns, they represent only a subset of the body. Some insights may generalize to other regions with similar roles, such as joints that govern ROM, supporting muscle groups, and related structures, but future work should extend our approach to lower-body and full-body exercises, where different biomechanical demands may surface new requirements for sensing and feedback design.
Second, in our PoC validation we relied on manual configuration of straps to set feedback intensity levels with six participants. While sufficient for this study, this approach introduced setup delays and required facilitator involvement. While we deliberately incorporated certain underexplored design dimensions into the PoC, there remains an opportunity for more comprehensive evaluation of other aspects, such as the semantics of various feedback modalities, especially in scenarios involving multiple simultaneous feedback cues. Coordinating such cues effectively to minimize cognitive load and maintain understandable semantics in time-sensitive settings represents an important next step toward supporting safe training. Additionally, higher system fidelity and a larger number of users may help move the vision closer to a usable product.
Third, our strap tension measurements (7–15 N across different locations) revealed an important limitation for wearable systems relying on passive, strap-based feedback. Stronger constraining cues require higher strap tensions, yet higher tensions make it more difficult to maintain comfort and stability without bulky motors or anchoring structures. Designing garments that can distribute forces evenly across the body while remaining flexible and wearable in dynamic exercise contexts remains an open challenge.
@@ 10. CONCLUSION
This paper presented FlexGuard, a design space for on-body wearable feedback to scaffold safety in strength training, derived from nine co-design workshops with trainer–trainee pairs using a technology probe that explored both extrinsic and intrinsic modalities. FlexGuard is articulated through dimensions of Sensing & Triggering and Feedback Delivery, providing designers with a framework to relate safety-relevant movement risks to appropriate forms of on-body feedback. We conceptually validated this design space through a speed dating study with seven beginners, which highlighted how users reason about feedback semantics, movement continuity, and adaptation across fast-changing training contexts. We further examined selected dimensions through a proof-of-concept prototype evaluated with six beginners, where results substantiated the relevance of the proposed dimensions and clarified assumptions about users’ underlying needs surfaced in the speed dating study. Together, these findings demonstrate that FlexGuard offers a valid and useful framework for designing on-body feedback systems that scaffold trainees’ awareness of and support for safety-relevant movement during strength training.